A car starts from rest and at the end of 80 seconds has a speed of 25 km/hr. What is the acceleration?
1
Expert's answer
2020-09-09T10:21:13-0400

Velocity for an accelerated motion changes as "v(t) = v_0 + a t", where "v_0" is initial velocity, "a" is acceleration and "t" is time.

In our case, "v_0 = 0" (start from rest), hence "v(t) = a t". From the last formula, "a = \frac{v(t)}{t} = \frac{25 \frac{km}{hr}{}}{80 s} = \frac{25 \frac{km}{hr}}{80/3600 hr} = 1125 \frac{km}{hr^2}"